1. Evaluate F(x)=3x^5 -5x^3 +2x^2 +6 for: 2x
2. Find the equation for the following line: A line with a slope of zero that passes through (2,-1)
3. Find the equation for the following line: A line with an undefined slope that passes through (-1,2)

Answer by stanbon(48516)

1. Evaluate F(x)=3x^5 -5x^3 +2x^2 +6 for: 2x
If you mean F(2x) you get:
F(2x) = 2(2x)^5 - 5(2x)^3 + 2(2x)^2 + 6
I'll let you work that out.
--------------------------------
2. Find the equation for the following line: A line with a slope of zero that passes through (2,-1)
If slope is zero the equation has the form y=k
But the point goes thru a point where y=-1
So the equation is y=-1
------------------------
3. Find the equation for the following line: A line with an undefined slope that passes through (-1,2)
If the slope is undefind the equation has the form x=k
But the point goes thru a point where x=1-
So the equation is x=-1
================
Cheers,
Stan H.